# Self-Guided Function Calling in Large Language Models via Stepwise Experience Recall

### v1

@@ Abstract
Sijia Cui1,2, Aiyao He1, Shuai Xu3, Hongming Zhang1, Yanna Wang1, Qingyang Zhang1, Yajing Wang4, Bo Xu 1 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Nanjing University of Information Science & Technology 4 Institute of Computing Technology, Chinese Academy of Sciences Correspondence: Yanna Wang<wangyanna2013@ia.ac.cn>, Bo Xu<boxu@ia.ac.cn>
@@ 1 Introduction
Large language models (LLMs) demonstrated remarkable capabilities through pretraining on large-scale corpora (Brown et al., 2020; Devlin et al., 2019; Touvron et al., 2023; Achiam et al., 2023; Denison et al., 2024; Team et al., 2023). However, due to the inherent limitations of neural network architecture, LLMs are unable to interact directly with the real world—an issue that cannot be resolved simply by scaling up the training data or model size. Function calling [^1] We use function calling and tool-use interchangeably in this paper. Qu et al. (2025); Qin et al. (2024a, b) serves as a fundamental mechanism that enables LLMs to interact with external systems. By invoking external tools, LLMs can integrate up-to-date knowledge and execute real-world tasks, thereby expanding the boundaries of LLM-based AI agents and driving advancements across various domains Hao et al. (2025); Theuma and Shareghi (2024); Zhong et al. (2023); Zhao et al. (2024).
@@
We introduce Stepwise ExperiencE Recall (SEER), a novel approach that enhances the multi-step tool-use capabilities of LLMs through fine-grained retrieval. It selects relevant trajectories by jointly considering task similarity, toolchain coverage, and intent alignment. SEER incrementally expands the experience pool by incorporating successful task trajectories, improving model performance over time. For tasks without explicit success signals, we adopt an LLM-as-a-judge mechanism Li et al. (2024) to assess task completion. This enables continuous online updates to the pool, allowing SEER to adapt to new tasks and evolving user demands. Our main contributions are:
@@ 2.1 Multi-step Function Calling
Recent progress in tool-augmented LLMs has focused on either freezing or training approaches Wang et al. (2024); Huang et al. (2023); Yu et al. (2024); Goldie et al. (2025). Many studies exploit LLMs’ in-context learning by prompting task descriptions and tool-use examples during inference Lu et al. (2023); Shen et al. (2023); Hsieh et al. (2023); Paranjape et al. (2023). For example, the ART framework Paranjape et al. (2023) retrieves similar multi-step reasoning and traces to guide models in generating intermediate steps and invoking functions. Other methods, such as StepTool Yu et al. (2024) and SWiRL Goldie et al. (2025), treat tool use as a reinforcement learning problem. StepTool applies step-wise reward shaping and policy gradients to improve decision-making based on tool success and task contribution. SWiRL generates synthetic multi-step tool-use data and uses step-wise RL with reward models to train without manual labels. In contrast, we introduce an experience replay approach that enhances multi-step tool use without additional training, using fine-grained replay to improve performance efficiently.
@@ 2.2 Self-improvement for LLM
Accelerated advancements in LLMs have intensified data scarcity issues, highlighting data bottlenecks as a major research challenge Villalobos et al. (2022). Self-improvement involving model-generated data such as feedback, instructions, and questions, has shown promise but often relies on heuristics and human validation for quality assurance Bai et al. (2022); Wang et al. (2022). Systems like ExpeL Zhao et al. (2024) leverage past task experiences to enhance decision-making at inference, and recent work by (Tian et al., 2024) integrates Monte Carlo Tree Search (MCTS) with language models, creating annotation-free self-improvement loops. However, these methods typically rely on static offline datasets, significantly limiting adaptability in practical scenarios. To address this, we propose an online updating experience pool that continuously supports in-context learning, enhancing inference quality and real-world adaptability.
@@ A.1 Intent Inference Prompt
The prompt construction process is showed in 4.
@@ A.2 Trajectory Evaluation Prompt
The prompt construction process is showed in 5.

### v2

@@ Abstract
Sijia Cui1,2, Aiyao He1, Shuai Xu3, Hongming Zhang1, Yanna Wang1, Qingyang Zhang1, Yajing Wang4, Bo Xu1 1The Key Laboratory of Cognition and Decision Intelligence for Complex Systems, Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Nanjing University of Information Science & Technology 4 Institute of Computing Technology, Chinese Academy of Sciences Correspondence: wangyanna2013@ia.ac.cn, boxu@ia.ac.cn
@@ 1 Introduction
Large language models (LLMs) demonstrated remarkable capabilities through pretraining on large-scale corpora (Brown et al., 2020; Devlin et al., 2019; Touvron et al., 2023; Achiam et al., 2023; Denison et al., 2024; Team et al., 2023). However, due to the inherent limitations of neural network architecture, LLMs are unable to interact directly with the real world—an issue that cannot be resolved simply by scaling up the training data or model size. Function calling [^1] We use function calling and tool-use interchangeably. Qu et al. (2025); Qin et al. (2024a, b) serves as a fundamental mechanism that enables LLMs to interact with external systems. By invoking external tools, LLMs can integrate up-to-date knowledge and execute real-world tasks, thereby expanding the boundaries of LLM-based AI agents and driving advancements across various domains Hao et al. (2025); Theuma and Shareghi (2024); Zhong et al. (2023); Zhao et al. (2024).
@@
We introduce Stepwise ExperiencE Recall (SEER) [^2] https://github.com/AI-Research-TeamX/SEER, a novel approach that enhances the multi-step tool-use capabilities of LLMs through fine-grained retrieval. It selects relevant trajectories by jointly considering task similarity, toolchain coverage, and intent alignment. SEER incrementally expands the experience pool by incorporating successful task trajectories, improving model performance over time. For tasks without explicit success signals, we adopt an LLM-as-a-judge mechanism Li et al. (2024) to assess task completion. This enables continuous online updates to the pool, allowing SEER to adapt to new tasks and evolving user demands. Our main contributions are:
@@ 2.1 Multi-step Function Calling
Recent progress in tool-augmented LLMs has focused on either freezing or training approaches Wang et al. (2024); Huang et al. (2023); Yu et al. (2024); Goldie et al. (2025). Many studies exploit LLMs’ in-context learning by prompting task descriptions and tool-use examples during inference Lu et al. (2023); Shen et al. (2023); Hsieh et al. (2023); Paranjape et al. (2023); Bai et al. (2024); Zhang et al. (2025); Yang et al. (2025); Xu et al. (2025); Cui et al. (2025). For example, the ART framework Paranjape et al. (2023) retrieves similar multi-step reasoning and traces to guide models in generating intermediate steps and invoking functions. Other methods, such as StepTool Yu et al. (2024) and SWiRL Goldie et al. (2025), treat tool use as a reinforcement learning problem Sutton and Barto (2018); Dong et al. (2020); Zhang and Yu (2020b). StepTool applies step-wise reward shaping and policy gradients to improve decision-making based on tool success and task contribution. SWiRL generates synthetic multi-step tool-use data and uses step-wise RL with reward models to train without manual labels. In contrast, we introduce an experience replay approach that enhances multi-step tool use without additional training, using fine-grained replay to improve performance efficiently.
@@ 2.2 Self-improvement for LLM
Accelerated advancements in LLMs have intensified data scarcity issues, highlighting data bottlenecks as a major research challenge Villalobos et al. (2022). Self-improvement involving model-generated data such as feedback, instructions, and questions, has shown promise but often relies on heuristics and human validation for quality assurance Bai et al. (2022); Wang et al. (2022). Systems like ExpeL Zhao et al. (2024) leverage past task experiences to enhance decision-making at inference, and recent work by (Tian et al., 2024) integrates Monte Carlo Tree Search (MCTS) Kocsis and Szepesvári (2006); Zhang and Yu (2020a) with language models, creating annotation-free self-improvement loops. However, these methods typically rely on static offline datasets, significantly limiting adaptability in practical scenarios. To address this, we propose an online updating experience pool that continuously supports in-context learning, enhancing inference quality and real-world adaptability.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Key R&D Program of China (No.2022ZD0116405).
Appendix
@@ A.1 Intent Inference Prompt
The prompt construction is shown in Figure 4.
@@ A.2 Trajectory Evaluation Prompt
The prompt construction is shown in Figure 5.